# Anisotropic examples of inflation-generating initial conditions for the big bang

### v1

@@ Abstract
The inflationary scenario, which states that the early universe underwent a brief but dramatic period of accelerated spatial expansion, has become the current paradigm of early universe cosmology. Although inflationary cosmology has its many successes, it does not (as of yet) have the status of an established physical theory. In this paper, we provide mathematical support for the inflationary scenario in a class of anisotropic spacetimes by generalizing the work in [19]. These anisotropic spacetimes satisfy certain initial conditions so that they are perfectly isotropic at the big bang but become less isotropic as time progresses. The resulting inflationary eras are a consequence of the energy-momentum tensor being dominated by a cosmological constant at the big bang.
@@ 1.1 Cosmic inflation
The inflationary scenario was proposed in the late 1970s and early 80s [11, 27, 16] as a solution to some problems in the standard big bang model, e.g., the flatness and horizon problems. It was soon realized that inflation can provide a framework for generating the seeds of the large-scale structures in our universe [22]. Observations of the anisotropies in the CMB performed by COBE, WMAP, and most recently by Planck[6] support these claims.
Given the many successes of the inflationary scenario, it is perhaps not too surprising that most papers on early universe cosmology give the impression that inflation has been firmly established and observationally proven. However there are many inflationary models that can be in agreement with observation [21]. In fact, any theory which predicts an almost flat universe with a nearly scale-invariant curvature power spectrum, small tensor-to-scalar ratio, and small Gaussian fluctuations would be in agreement with current data.
@@ 1.3 Geometrical initial conditions for the big bang
An origin point $\mathcal{O}$ for $u$ is the first main assumption in Theorem 2.4. The other main assumption is that the energy-momentum tensor $T$ approaches that of a perfect fluid at $\mathcal{O}$. See Definition 2.2. This assumption is more physically convincing than assuming that $T$ is exactly a perfect fluid (as in the FLRW models) since we expect small deviations from perfect isotropy in our universe. As Penrose reasons [25], the deviation from perfect isotropy in our universe is induced by gravitational collapse. Therefore it stands to reason that the universe becomes more isotropic towards the past and perhaps perfectly isotropic at the big bang. This is the crux of Definition 2.2.
@@ 2 The main theorem
    $\widetilde{\gamma}^{\prime}(0)$ exists and $\widetilde{\gamma}^{\prime}(0)=\lim_{t\to 0}\gamma^{\prime}(t)$,
where $\widetilde{\gamma}\colon[0,b)\to M_{\text{{\rm ext}}}$ is the extended curve defined by $\widetilde{\gamma}(0)=\mathcal{O}$ and $\widetilde{\gamma}(t)=\gamma(t)$ for $t>0$. Continuity of the metric implies $\widetilde{\gamma}^{\prime}(0)$ is a unit future directed timelike vector.
Remarks. Definition 2.1 is supposed to model the behavior of the comoving observers in Figure 1 (right). It is essentially the same as assumption (b) in [19, Thm. 2.2]. Actually, Definition 2.1 is slightly stronger; we assume this stronger assumption since it’s easier to state and all the examples in section 3 will satisfy it.
@@
Remark. Definition 2.2 relaxes the requirement that $T$ is identically a perfect fluid in assumption (a) of [19, Thm. 2.2]. Moreover, its more physically convincing: the universe is not perfectly isotropic right now, but observational evidence suggests that the universe becomes more isotropic as we go further back in time towards the big bang.
@@
Remarks. In [19, Thm. 2.2], it was assumed that the manifold $M$ satisfies $M=I^{+}(\mathcal{O},M_{\text{{\rm ext}}})$. Definition 2.3 relaxes this requirement and is a much weaker assumption. It will hold for the examples constructed in section 3. Also, conditions (i) and (ii) in Definition 2.3 will always be satisfied by continuity of the metric and applying the Gram-Schmidt orthogonalization process appropriately. The heart of Definition 2.3 is condition (iii) and is the motivation for the terminology “timelike convex near $\mathcal{O}$.”
We are now ready to state our main theorem which generalizes [19, Thm. 2.2].
@@
By assumption (b), $T_{\rm perfect}-T$ extends continuously to $M\cup\{\mathcal{O}\}$, and its continuous extension is zero at $\mathcal{O}$. Also $T$ extends continuously to $M\cup\{\mathcal{O}\}$ by assumption (d). Therefore $u_{*}\otimes u_{*}$ extends continuously to $M\cup\{\mathcal{O}\}$. As in the proof of [19, Thm. 2.2], this implies that the vector field $u$ extends continuously to $M\cup\{\mathcal{O}\}$. However, assumptions (c) and (e) prove that $u$ does not extend continuously. Heuristically, this can be seen in Figure 1 (right). Rigorously, this follows from an analogous contradiction argument used in the proof of [19, Thm. 2.2]. Thus we have $\rho_{0}=-p_{0}$.
@@ 3.1 The examples
for some arbitrary smooth [^4] More generally $\Phi(\zeta)$ only needs to be a $C^{2}$ function on a neighborhood of $\zeta=0$. For simplicity we assume $\Phi(\zeta)$ is smooth on all of $\mathbb{R}$. function $\Phi(\zeta)$ on $\mathbb{R}$. Here $\zeta=\zeta(t,r)$ is given by
@@
The spacetime extension $(M_{\text{{\rm ext}}},g_{\text{{\rm ext}}})$ of $(M_{\varepsilon},g)$ is simply defined by extending $g$ to all $\mathbb{R}^{4}\approx M_{\text{{\rm ext}}}$. In fact the metric is $C^{2}$ on $M_{\text{{\rm ext}}}$, which follows from the assumptions on $f(r)$.
@@
The unit future directed timelike vector field $u$ (whose integral curves will be the comoving observers) will be given by normalized gradient of $\zeta$:
@@
$R_{\alpha\beta}=-2(\operatorname{Hess}\Phi)_{\alpha\beta}+2\nabla_{\alpha}\Phi\nabla_{\beta}\Phi-\big{(}\Delta\Phi+2|d\Phi|_{\eta}^{2}\big{)}\eta_{\alpha\beta}.$ (3.8)
@@
$=-2(\operatorname{Hess}\Phi)_{\alpha\beta}+2\nabla_{\alpha}\Phi\nabla_{\beta}\Phi+\Delta\Phi\,\eta_{\alpha\beta}+2|d\Phi|^{2}_{\eta}\,\eta_{\alpha\beta},$
@@
Remark. We emphasize that we have applied Theorem 2.4 to the spacetime $M_{\varepsilon}$ and not $M$, but this is sufficient for our analysis in the next section.
@@ 3.2 Existence of inflationary eras in the examples
The assumption in (3.15) is what we mean by “the cosmological constant appears as an initial condition.” It holds for a class of Milne-like spacetimes, see [19, eq. (1.11)]. In fact, our main result, Theorem 2.4, is essentially an anisotropoic generalization of this.
@@
Recall $u$ is given by (3.4). By construction $u$ is orthogonal to the spacelike hypersurfaces of constant $\zeta$. In Figure 1 (right), one should image that the spacelike hypersurfaces $\tau=\rm{constant}$ are replaced with $\zeta=\rm{constant}$. In the terminology of [24, p. 359], $u$ is “synchronizable,” but it is not necessarily “proper time synchronizable.” The latter occurs if and only if $u$ is geodesic which occurs if and only if $f(r)=r$, see eq. (3.23). (Recall $f(r)=r$ corresponds to a Milne-like spacetime, and we know $u$ is geodesic in this case.)
@@
Along each comoving observer $\gamma(\tau)$, we define a generalized scale factor $\mathfrak{a}(\tau)$ by [^6] The generalized scale factor $\mathfrak{a}(\tau)$ is also known as an “average length scale”, see [7, eq. (2.14)]. In the isotropic FLRW setting, $\mathfrak{a}(\tau)$ is simply the scale factor $a(\tau)$, and the mean curvature $H$ corresponds to the Hubble parameter. It’s a happy coincidence that the notation for the mean curvature and the Hubble parameter happen to coincide.
@@
We have $H(\tau)\approx\frac{1}{\tau}$ for $\tau$ small along each comoving observer $\gamma(\tau)$, see equation (3.21). Since $\mathfrak{a}(\tau)=\exp(\int_{\tau_{0}}^{\tau}H)$ for some arbitrary time $\tau_{0}$, it follows that
@@
For FLRW spacetimes, Friedmann’s second equation (3.14) is used to analyze the acceleration of the scale factor. In the anisotropic setting, the generalization of Friedmann’s second equation is the Raychaudhuri equation [13, eq. (4.26)],
@@
Our goal is to compute all the terms on the right-hand side of (3.19) for points along a comoving observer near $\mathcal{O}$. First, using the conclusions of Theorem 2.4 and eq. (3.13), sufficiently close to the origin $\mathcal{O}$, we have
@@
$2\sigma^{2}\,=\,\frac{2e^{-2\Phi(\zeta)}}{3[t^{2}f^{\prime}(r)^{2}-f(r)^{2}]}\left[f^{\prime}(r)-\frac{f(r)}{r}-\frac{t^{2}f(r)f^{\prime}(r)f^{\prime\prime}(r)}{t^{2}f^{\prime}(r)^{2}-f(r)^{2}}\right]^{2}.$
@@
In other words $2\sigma^{2}$ extends continuously to $M_{\varepsilon}\cup\{\mathcal{O}\}$ and takes on the value 0 at $\mathcal{O}$.
@@ 3.3 Remarks on proving anisotropy
First, the definition of an “isotropic spacetime” is not consistent throughout the literature. See [2] and [26] and references therein. We will adopt the definition in [24, Ch. 12] since, as discussed in [2], this definition is the optimal one as it implies that the spacetime is isometric to a subset of an FLRW spacetime, see [24, Prop. 12.6] and [2, Thm. 2.1].
Therefore any spacetime that is not isometric to a subset of an FLRW model is anisotropic according to [24]. For the examples constructed in section 3.1, if $f(r)=r$ then they are isometric to a subclass of Milne-like spacetimes which are a subclass of $k=-1$ FLRW models, and hence they are isotropic. Moreover, regardless of the form of $f(r)$, if the conformal factor is identically $1$, then the spacetime is isometric to a subset of Minkowski spacetime which is clearly isotropic. (However, in this case, the vector field defining the comoving observers changes.) This suggests that if $f(r)$ is not identically $r$ and the conformal factor is not constant, then the resulting spacetime is not a subset of an FLRW spacetime and hence is anisotropic. We believe such a statement can be proven rigorously. However, in this section, we will content ourselves with the following algorithm: Pick functions $f(r)$ and $\Phi(\zeta)$. The steps below show how to verify that the corresponding spacetime $(M,g)$ from section 3.1 is anisotropic.
Seeking a contradiction, suppose $(M,g)$ is in fact isometric to a subset of an FLRW spacetime. Since FLRW spacetimes satisfy the Einstein equations with a perfect fluid [24, Thm. 12.11], there is a unit future directed timelike vector field $\widetilde{u}$ on $M$ such that $T$ is a perfect fluid with respect to $\widetilde{u}$. There exist functions $a$, $b$, $c$, $d$ such that
@@ 4 Summary and outlook
We believe that differential geometry (and geometric analysis in particular) has a role to play in the investigation of initial conditions for the big bang. The work presented in this paper should be thought of as a “proof of concept” of this proposal. Our work can be generalized in many ways. In particular, although our examples are not necessarily isotropic, they are still spherically symmetric. So a natural generalization is to reproduce the analysis in sections 3.1 and 3.2 with non-spherically symmetric spacetimes. Also, our examples are anisotropic versions of $k=-1$ FLRW spacetimes. What about $k=0$ FLRW spacetimes? In this case one would want to apply [10, Thm. 5.2] or a suitable generalization thereof. Lastly, it remains to be seen if the results in [9] can be used to generate comoving observers with an origin point $\mathcal{O}$.

### v2

@@ Abstract
The inflationary scenario, which states that the early universe underwent a brief but dramatic period of accelerated spatial expansion, has become the current paradigm of early universe cosmology. Although inflationary cosmology has its many successes, it does not (as of yet) have the status of an established physical theory. In this paper, we provide mathematical support for the inflationary scenario in a class of anisotropic spacetimes by generalizing the work in [21]. These anisotropic spacetimes satisfy certain initial conditions so that they are perfectly isotropic at the big bang but become less isotropic as time progresses. The resulting inflationary eras are a consequence of the initial conditions which force the energy-momentum tensor to be dominated by a cosmological constant at the big bang.
@@ 1.1 Cosmic inflation
The inflationary scenario was proposed in the late 1970s and early 80s [13, 29, 18] as a solution to some problems in the standard big bang model, e.g., the flatness and horizon problems. It was soon realized that inflation can provide a framework for generating the seeds of the large-scale structures in our universe [24]. Observations of the anisotropies in the CMB radiation performed by COBE, WMAP, and most recently by Planck[8] support these claims.
Given the many successes of the inflationary scenario, it is perhaps not too surprising that most papers on early universe cosmology give the impression that inflation has been firmly established and observationally proven. However there are many inflationary models that can be in agreement with observation [23]. In fact, any theory which predicts an almost flat universe with a nearly scale-invariant curvature power spectrum, small tensor-to-scalar ratio, and small Gaussian fluctuations would be in agreement with current data, e.g., [5].
@@ 1.3 Geometrical initial conditions for the big bang
An origin point $\mathcal{O}$ for $u$ is the first main assumption in Theorem 2.4. The other main assumption is that the energy-momentum tensor $T$ approaches that of a perfect fluid at $\mathcal{O}$. See Definition 2.2. This assumption is more physically convincing than assuming that $T$ is exactly a perfect fluid (as in the FLRW models) since we expect small deviations from perfect isotropy in our universe. Therefore the perspective taken here is that the universe began in a state of perfect isotropy at the big bang. This is the crux of Definition 2.2. Moreover, this perspective is reinforced in our examples since the shear vanishes towards the big bang, see eq. (3.23).
@@ 2 The main theorem
    $\widetilde{\gamma}^{\prime}(0)$ exists and $\widetilde{\gamma}^{\prime}(0)=\displaystyle\lim_{\tau\to 0}\gamma^{\prime}(\tau)$,
where $\widetilde{\gamma}\colon[0,b)\to M_{\text{{\rm ext}}}$ is the extended curve defined by $\widetilde{\gamma}(0)=\mathcal{O}$ and $\widetilde{\gamma}(\tau)=\gamma(\tau)$ for $\tau>0$. Continuity of the metric implies $\widetilde{\gamma}^{\prime}(0)$ is a unit future directed timelike vector.
Remarks. Definition 2.1 is supposed to model the behavior of the comoving observers in Figure 1 (right). It is essentially the same as assumption (b) in [21, Thm. 2.2]. Actually, Definition 2.1 is slightly stronger; we assume this stronger assumption since it’s easier to state and all the examples in section 3 will satisfy it.
@@
Remark. Definition 2.2 relaxes the requirement that $T$ is identically a perfect fluid in assumption (a) of [21, Thm. 2.2]. Moreover, it’s more physically convincing: FLRW models have perfect fluid energy-momentum tensors, and we expect that an FLRW model approximates our universe better as we go back in time towards the big bang.
@@
Remarks. In [21, Thm. 2.2], it was assumed that the manifold $M$ satisfies $M=I^{+}(\mathcal{O},M_{\text{{\rm ext}}})$. Definition 2.3 relaxes this requirement and is a much weaker assumption. It will hold for the examples constructed in section 3. Also, conditions (i) and (ii) in Definition 2.3 will always be satisfied by continuity of the metric and applying the Gram-Schmidt orthogonalization process appropriately. The heart of Definition 2.3 is condition (iii) and is the motivation for the terminology “timelike convex near $\mathcal{O}$.”
We are now ready to state our main theorem which generalizes [21, Thm. 2.2].
@@
By assumption (b), $T_{\rm perfect}-T$ extends continuously to $M\cup\{\mathcal{O}\}$, and its continuous extension is zero at $\mathcal{O}$. Also $T$ extends continuously to $M\cup\{\mathcal{O}\}$ by assumption (d). Therefore $u_{*}\otimes u_{*}$ extends continuously to $M\cup\{\mathcal{O}\}$. As in the proof of [21, Thm. 2.2], this implies that the vector field $u$ extends continuously to $M\cup\{\mathcal{O}\}$. However, assumptions (c) and (e) prove that $u$ does not extend continuously. Heuristically, this can be seen in Figure 1 (right). Rigorously, this follows from an analogous contradiction argument used in the proof of [21, Thm. 2.2]. Thus we have $\rho_{0}=-p_{0}$.
@@ 3.1 The examples
for some arbitrary smooth [^4] More generally $\Phi(\zeta)$ only needs to be a $C^{2}$ function on a neighborhood of $\zeta=0$, but, for simplicity, we assume $\Phi(\zeta)$ is smooth on all of $\mathbb{R}$. function $\Phi(\zeta)$ on $\mathbb{R}$. Here $\zeta=\zeta(t,r)$ is given by
@@
The spacetime extension $(M_{\text{{\rm ext}}},g_{\text{{\rm ext}}})$ of $(M,g)$ is simply defined by extending $g$ to all $\mathbb{R}^{4}\approx M_{\text{{\rm ext}}}$. In fact the metric is $C^{2}$ on $M_{\text{{\rm ext}}}$, which follows from the assumptions on $f(r)$.
@@
$\mathcal{O}$$t$$x^{i}$$\tau=\text{constant}$ $t=\mu r$$M:=\{(t,x,y,z)\mid t>r\},\quad g=e^{2\Phi(\tau(t,r))}\eta$with $\tau(t,r)=\frac{t^{2}}{2}-\frac{r^{2}}{2}.$
[…]
Figure 2: On the left is a Milne-like spacetime represented in conformal Minkowskian coordinates. On the right, the anisotropic examples constructed in this section. They are constructed to look like a Milne-like spacetime around the origin $\mathcal{O}$. The comoving observers (i.e., the integral curves of $u$) still emanate from the origin, and the manifold is still foliated by slices orthogonal to the comoving observers.
The unit future directed timelike vector field $u$ (whose integral curves are the comoving observers) will be given by normalized gradient of $\zeta$:
@@
$R_{\alpha\beta}=-2(\operatorname{Hess}\Phi)_{\alpha\beta}+2\nabla_{\alpha}\Phi\nabla_{\beta}\Phi-\big{(}\square\Phi+2|d\Phi|_{\eta}^{2}\big{)}\eta_{\alpha\beta}.$ (3.8)
@@
$=-2(\operatorname{Hess}\Phi)_{\alpha\beta}+2\nabla_{\alpha}\Phi\nabla_{\beta}\Phi+\square\Phi\,\eta_{\alpha\beta}+2|d\Phi|^{2}_{\eta}\,\eta_{\alpha\beta},$
@@
Remark. We emphasize that we have applied Theorem 2.4 to the spacetime $M_{\varepsilon}$ and not $M$; this is sufficient for the analysis in the next section.
@@ 3.2 Existence of inflationary eras in the examples
The assumption in (3.15) is what we mean by “the cosmological constant appears as an initial condition.” It holds for a class of Milne-like spacetimes, see [21, eq. (1.11)]. In fact, our main result, Theorem 2.4, is essentially an anisotropoic generalization of this.
@@
$t$$x^{i}$$t=\mu f(r)$$\mathcal{O}$
Figure 3: We prove the existence of inflationary eras along a fixed comoving observer by computing the terms on the right-hand side of the Raychaudhuri equation (3.19) in a small neighborhood in $M$ about the origin.
Recall $u$ is given by (3.4). By construction $u$ is orthogonal to the spacelike hypersurfaces of constant $\zeta$. In Figure 1 (right), one should image that the spacelike hypersurfaces $\tau=\rm{constant}$ are replaced with $\zeta=\rm{constant}$. In the terminology of [26, p. 359], $u$ is “synchronizable,” but it is not necessarily “proper time synchronizable.” The latter occurs if and only if $u$ is geodesic which occurs if and only if $f(r)=r$, see eq. (3.24). (Recall $f(r)=r$ corresponds to a Milne-like spacetime, and we know $u$ is geodesic in this case.)
@@
Along each comoving observer $\gamma(\tau)$, we define a generalized scale factor $\mathfrak{a}(\tau)$ by [^6] The generalized scale factor $\mathfrak{a}(\tau)$ is also known as an “average length scale”, see [9, eq. (2.14)]. In the isotropic FLRW setting, $\mathfrak{a}(\tau)$ is simply the scale factor $a(\tau)$, and the mean curvature $H$ corresponds to the Hubble parameter. It’s a happy coincidence that the notation for the mean curvature and the Hubble parameter happen to coincide.
@@
We have $H(\tau)\approx\frac{1}{\tau}$ for $\tau$ small along each comoving observer $\gamma(\tau)$, see eq. (3.21) below. Since $\mathfrak{a}(\tau)=\exp(\int_{\tau_{0}}^{\tau}H)$ for some arbitrary time $\tau_{0}$, it follows that
@@
For FLRW spacetimes, Friedmann’s second equation (3.14) is used to analyze the acceleration of the scale factor. In the anisotropic setting, the generalization of Friedmann’s second equation is the Raychaudhuri equation [15, eq. (4.26)],
@@
Our goal is to compute all the terms on the right-hand side of (3.19) for points along a comoving observer near $\mathcal{O}$ (see Figure 3). First, using the conclusions of Theorem 2.4 and eq. (3.13), sufficiently close to the origin $\mathcal{O}$, we have
@@
$2\sigma^{2}\,=\,\frac{2e^{-2\Phi(\zeta)}}{3[t^{2}f^{\prime}(r)^{2}-f(r)^{2}]}\left[f^{\prime}(r)-\frac{f(r)}{r}-\frac{t^{2}f(r)f^{\prime}(r)f^{\prime\prime}(r)}{t^{2}f^{\prime}(r)^{2}-f(r)^{2}}\right]^{2}.$ (3.22)
@@
In other words $2\sigma^{2}$ extends continuously to $M_{\varepsilon}\cup\{\mathcal{O}\}$ and takes on the value 0 at $\mathcal{O}$. Geometrically, this “isotropization” effect is a consequence of the $u$-orthogonal hypersurfaces becoming more hyperbolic as we approach the origin $\mathcal{O}$.
@@ 3.3 Remarks on proving anisotropy
First, the definition of an “isotropic spacetime” is not consistent throughout the literature. See [2] and [28] and references therein. We will adopt the definition in [26, Ch. 12] since, as discussed in [2], this definition is the optimal one as it implies that the spacetime is isometric to a subset of an FLRW spacetime, see [26, Prop. 12.6] and [2, Thm. 2.1].
Therefore any spacetime that is not isometric to a subset of an FLRW model is anisotropic according to [26]. For the examples constructed in section 3.1, if $f(r)=r$ then they are isometric to a subclass of Milne-like spacetimes which are a subclass of $k=-1$ FLRW models, and hence they are isotropic. Moreover, regardless of the form of $f(r)$, if the conformal factor is identically $1$, then the spacetime is isometric to a subset of Minkowski spacetime which is clearly isotropic. (This shows that it is not sufficient to simply recognize that the shear term (3.22) is nonzero. However, in this case, the vector field defining the comoving observers changes.) This suggests that if $f(r)$ is not identically $r$ and the conformal factor is not constant, then the resulting spacetime is not a subset of an FLRW spacetime and hence is anisotropic. We believe such a statement can be proven rigorously. However, in this section, we will content ourselves with the following algorithm: Pick functions $f(r)$ and $\Phi(\zeta)$. The steps below show how to verify that the corresponding spacetime $(M,g)$ from section 3.1 is anisotropic.
Seeking a contradiction, suppose $(M,g)$ is in fact isometric to a subset of an FLRW spacetime. Since FLRW spacetimes satisfy the Einstein equations with a perfect fluid [26, Thm. 12.11], there is a unit future directed timelike vector field $\widetilde{u}$ on $M$ such that $T$ is a perfect fluid with respect to $\widetilde{u}$. There exist functions $a$, $b$, $c$, $d$ such that
@@ 4 Summary and outlook
Our examples exhibit isotropization towards the past and, in fact, are perfectly isotropic at the big bang $\mathcal{O}$, see (3.23). Our isotropization-towards-the-past result is consistent with a universe starting from special initial conditions. This is unlike results related to the cosmic no-hair conjecture (see Wald’s original paper [30] or some more recent work, e.g., [3]), where isotropization occurs towards the future.
@@
We believe that differential geometry (and geometric analysis in particular) has a role to play in the investigation of initial conditions for the big bang. The work presented in this paper should be thought of as a “proof of concept” of this proposal. Our work can be generalized in many ways. In particular, although our examples are not necessarily isotropic, they are still spherically symmetric. So a natural generalization is to reproduce the analysis in sections 3.1 and 3.2 with non-spherically symmetric spacetimes. Also, our examples are anisotropic versions of $k=-1$ FLRW spacetimes. What about $k=0$ FLRW spacetimes? In this case one would want to apply [12, Thm. 5.2] or a suitable generalization thereof. Lastly, it remains to be seen if the results in [11] can be used to generate comoving observers with an origin point $\mathcal{O}$.